# Discount Model Search for Quality Diversity Optimization in High-Dimensional Measure Spaces

### v1

@@ Appendix A Results
A.1. Comparability of Results to Prior Work
Variations of the linear projection and arm repertoire domains have appeared in a number of prior works, each with slightly different setups than the one in our work. Here, to facilitate better comparability with the results in prior work, we clarify the distinctions between our setup and previous ones. The first difference is in the order of magnitude of the scores reported in our work. Prior work like Fontaine 2021 (Fontaine and Nikolaidis, 2021) and Fontaine 2023 (Fontaine and Nikolaidis, 2023) reported normalized QD scores, which are QD scores divided by the number of cells in the archive (10,000). Those works also normalized objective values to be 0-100, whereas we normalize to 0-1 to make the objective values more amenable to the discount model. These two changes explain the difference in orders of magnitude between our results and those works: when divided by 10,000 and multiplied by 100, our results for CMA-MAE, MAP-Elites, and MAP-Elites (line) on 2D LP (Sphere), 2D LP (Rastrigin), and Arm Repertoire are nearly identical to those for LP (sphere), LP (Rastrigin), and Arm Repertoire in Fontaine 2023 (our remaining benchmark domains were not included in Fontaine 2023).
@@ A.1. Comparability of Results to Prior Work
Figure 3. Mean and standard error of the mean for QD Score and Coverage of each algorithm in each domain. Standard error may not be visible in some plots.
@@
Figure 4. Mean and standard error of the mean for QD Score and Coverage of each algorithm in each domain. Standard error may not be visible in some plots.
@@
Figure 5. A random subset of images generated by DMS in the TA (MNIST) domain, where desired measures are sampled from the MNIST dataset. The goal in this domain is to arrange triangles to look like the given MNIST digits. Each rendered triangle image is shown to the left of its corresponding MNIST digit.
Figure 6. A random subset of images generated by DMS in the TA (F-MNIST) domain, where desired measures are sampled from the Fashion MNIST dataset. The goal in this domain is to arrange triangles to look like the images of fashion items. Each rendered triangle image is shown to the left of its corresponding fashion item.
Figure 7. Heatmaps of a randomly selected archive produced by each algorithm in domains with 2D measure spaces. Each row contains heatmaps for a single domain. The axes of the heatmaps are the measures, while the color of each cell indicates the objective value. Notably, the heatmaps show how DMS achieves high coverage of the measure space. They also show how DDS achieves good coverage but cannot achieve high objective values since it is a diversity optimization algorithm.
@@
Figure 8. Progression of the archive and discount model in DMS in the 2D LP (Sphere) benchmark. The left heatmap shows the archive, while the right heatmap shows the discount model. To plot the discount model, we computed its output at points in a $200\times 200$ grid in measure space. The discount model heatmap also shows the dataset $\mathcal{D}_{A}$ of points on a given iteration — blue circles indicates points created with solutions from the emitters, and yellow triangles indicate empty points. On Iteration 0, the discount model initializes to output $f_{min}$ everywhere. On Iteration 250, as the emitters begin to populate the archive, the discount model begins to output higher values in areas that have been explored. However, unexplored areas still maintain low values (shown as dark colors) due to the empty points in the dataset. On further iterations, the discount model outputs higher and higher values as the emitters populate the archive further, until it outputs high values nearly everywhere on Iteration 10000.
Table 2. Welch’s one-way ANOVA results in each domain. All $p$-values are less than 0.001.
@@
Table 3. Pairwise comparisons (Games-Howell test) of the QD Score of each algorithm.
@@
Table 4. Pairwise comparisons (Games-Howell test) of the Coverage of each algorithm.
@@
Table 5. Computation (wallclock) time (in seconds) of each algorithm in each domain. We show the mean and standard error of the mean over 20 trials for the benchmark domains and 5 trials for the QDDM domains.
2D LP (Sphere) 10D LP (Sphere) 20D LP (Sphere) 50D LP (Sphere)
DMS 397.83 ±0.37 876.33 ±44.22 2,797.61 ±6.57 7,843.98 ±228.28
CMA-MAE 121.13 ±0.10 333.92 ±1.70 1,316.87 ±22.32 4,707.92 ±86.53
DDS 1,918.68 ±22.42 3,535.78 ±1.53 1,481.18 ±30.98 2,600.24 ±78.20
MAP-Elites 39.01 ±0.21 190.25 ±2.46 873.06 ±0.35 4,523.16 ±51.12
2D LP (Rastrigin) 10D LP (Rastrigin) 2D LP (Flat) 10D LP (Flat)
DMS 707.20 ±21.31 711.08 ±30.76 683.53 ±27.73 723.34 ±32.31
CMA-MAE 182.07 ±0.36 465.80 ±3.81 165.33 ±0.25 800.12 ±0.92
DDS 1,957.29 ±0.87 3,598.86 ±2.56 1,937.41 ±1.51 3,581.79 ±2.05
MAP-Elites (line) 39.74 ±0.14 274.86 ±0.62 30.84 ±0.09 339.73 ±0.42
MAP-Elites 51.27 ±0.17 367.56 ±1.57 38.41 ±0.14 451.46 ±0.41
Arm Repertoire TA (MNIST) TA (F-MNIST) LSI (Hiker)
DMS 687.85 ±27.18 489.90 ±12.55 546.60 ±21.46 4,740.24 ±482.74
CMA-MAE 130.83 ±0.26 495.95 ±13.12 525.42 ±19.22 4,690.12 ±654.47
DDS 1,577.54 ±1.70
MAP-Elites 32.80 ±0.08 291.98 ±1.46 309.01 ±3.92 3,460.71 ±5.86
Appendix B Hyperparameters
Table 6. Hyperparameters.
@@ Appendix B Hyperparameters
Table 6 lists the hyperparameters of DMS and all baseline algorithms. All algorithms run for 10,000 iterations. The number of solutions generated and evaluated on each iteration is equal across all algorithms. In DMS, CMA-MAE, and DDS, this number is equivalent to the number of emitters $W$ times the emitter batch size $\lambda$. In the two MAP-Elites algorithms, this number is equivalent to the batch size $\lambda$. For DDS, we use the KDE version (“DDS-KDE”) (Lee et al., 2024). In all domains, the objective is normalized to be between 0 and 1, so the minimum objective $f_{min}$ is set to 0. For the benchmark domains, parameters for the baselines are adapted from prior work (Fontaine and Nikolaidis, 2023; Lee et al., 2024).
@@
Restart Rule. The restart rule refers to the conditions upon which emitters are restarted from solutions in the archive. “No Imp.” refers to restarting when the emitter no longer discovers solutions that are added to the archive (i.e., solutions that improve the archive) (Fontaine et al., 2020). “Basic” refers to restarting only when default CMA-ES (Hansen, 2016) termination rules are met. An integer value $R$ refers to restarting every $R$ iterations (Fontaine and Nikolaidis, 2023). We study the role of the restart rule in DMS in further detail in Appendix C.
@@
B.1. On the Choice of MLP Discount Models
@@
B.2. On Nearest Neighbor Search in Archives
@@
Appendix C Ablation Study
C.1. Archive Learning Rate
@@
C.2. Empty Points
C.3. Restart Rule
The “restart rule” in DMS and CMA-MAE refers to the conditions upon which the emitters restart search from a new emitter in the archive (Algorithm 1, line 1-1). By default, emitters in DMS and CMA-MAE use a “basic” restart rule, which makes the emitter restart when the default termination conditions for CMA-ES (Hansen, 2016) are met, e.g., the area of the search distribution becomes too small. While tuning DMS, we found it helpful in some domains to make the emitters restart on a fixed schedule as introduced in prior work (Fontaine and Nikolaidis, 2023), e.g., restarting every 100 iterations. To better understand the effect of the restart rule in DMS, we thus present an ablation where we run DMS in the benchmark domains with both a “basic” restart rule and a restart rule of 100.
Figure 9. Mean and standard error of the mean of QD Score and Coverage when varying the archive learning rate $\alpha$ in DMS in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@ C.3. Restart Rule
Figure 10. Mean and standard error of the mean of QD Score and Coverage when varying the number of empty points $n_{empty}$ in DMS in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@
Iteration 0
Iteration 250
Iteration 10000
Figure 11. Similar to Fig. 8, this figure shows how the archive and discount model in DMS progress across iterations. However, this time, DMS does not train the discount model with any empty points, i.e., $n_{empty}=0$. As a result, the discount model takes on arbitrary values in areas of the measure space that have not been explored yet, as evinced by the high values across the discount model heatmap on Iteration 250 and 10000. Because the discount values are high everywhere, the emitters in DMS mistakenly believe they have explored all areas of the measure space, even though the archive is essentially empty.
Table 7. Mean and standard error of the mean of QD Score and Coverage with a basic restart rule and a restart rule of 100 in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@
Appendix D Domain Details
D.1. Linear Projection
@@
Figure 12. Irwin-Hall distribution when summing $r$ uniformly random variables with range [5.12,5.12][-5.12,5.12].
@@
D.2. Arm Repertoire
Arm Repertoire (Cully and Demiris, 2017; Vassiliades and Mouret, 2018) considers solutions ${\bm{\theta}}\in\mathbb{R}^{n}$ that represent the $n$ joint angles of a planar robotic arm with $n$ joints and $n$ links of length 1. The measure function, ${\bm{m}}({\bm{\theta}})$, uses forward kinematics to compute the $(x,y)$ position of the end effector — thus, the measure space is two-dimensional. The objective function seeks to minimize the variance between each joint angle:
@@ D.2. Arm Repertoire
D.3. Triangle Arrangement (TA)
This QDDM domain considers arranging triangles to create images, as inspired by prior work (Tian and Ha, 2022). We consider grayscale images in our work, with the triangles all drawn on a black background. Each triangle is parameterized by the following eight values:
@@ D.3. Triangle Arrangement (TA)
D.4. Latent Space Illumination (LSI)
We construct this QDDM domain by adapting prior work that introduced LSI domains with 2D measures (Fontaine and Nikolaidis, 2021, 2023), and prior work that shows how to guide StyleGAN3 to generate images that match text prompts (Herrera-Berg, 2021). In our domain, which we refer to as LSI (Hiker), the goal is to generate face images of hikers in different landscapes, e.g., a hiker who is ready for the mountains.
@@ Appendix E Implementation
Compute Usage. In benchmark domains, we run all 20 trials of each algorithm in parallel, which for CMA-MAE and the two MAP-Elites variants takes about 5 minutes for domains with 2D measure spaces and 15 minutes for domains with 10D measure spaces. DDS takes 30 minutes in domains with 2D measure spaces and 1 hour in domains with 10D measure spaces. DMS takes 15 minutes in domains with 2D measure spaces and 30 minutes in domains with 10D measure spaces. In QDDM domains, we run algorithms serially. Each run in TA (MNIST) and TA (F-MNIST) takes 10 minutes, while a run in LSI (Hiker) takes 1 hour. For more precise wallclock times, refer to Table 5. In all domains, we train the discount model of DMS on the GPU. The discount model typically requires 300-400MB of GPU memory. In TA (MNIST) and TA (F-MNIST), we accelerate rendering of the triangles on the GPU. In LSI (Hiker), we accelerate StyleGAN3 and CLIP with the GPU. The final experimental results required 100GB of storage.
@@
Appendix F Differentiable Quality Diversity in QDDM Domains
While our work considers a black-box QD setting as defined in Sec. 2, prior work (Fontaine and Nikolaidis, 2021) introduced differentiable quality diversity (DQD), a QD setting where the objective and measure functions are first-order differentiable. Since DQD methods make assumptions that DMS and our other baselines do not, viz., that the objective and measures are differentiable, we have not included DQD methods in our experiments. Nevertheless, here we discuss considerations in applying DQD methods to QDDM domains.
The primary consideration is that DQD methods, such as CMA-MEGA (Fontaine and Nikolaidis, 2021), were designed with the assumption of a high-dimensional solution space that could be reduced to a low-dimensional objective-measure space, under the assumption that the measure space is low-dimensional. To elaborate, during operation, CMA-MEGA and other DQD methods compute a gradient for the objective and for each measure. In low-dimensional measure spaces, e.g., a 2D measure space, this works well because only a few gradients, 1 for the objective and 2 for the measures in this case, need to be computed. In contrast, in QDDM domains, the measure spaces are much higher-dimensional than even the solution space. In LSI (Hiker), the solution space is 512D while the measure space is 196,608D ($256\times 256\times 3$ RGB image). In the TA domains, the solution space is 240D while the measure space is 784D (it is also worth noting that the TA domains are non-differentiable due to the rendering process). As such, running DQD in these domains would mean computing up to 196,608 measure gradients on every iteration, which is computationally intractable. Given this limitation, we believe developing a DQD method that operates in QDDM domains is an exciting avenue for future work.

### v2

@@ Appendix B Results
B.1 Comparability of Results to Prior Work
Variations of the linear projection and arm repertoire domains have appeared in a number of prior works, each with slightly different setups than the one in our work. Here, to facilitate better comparability with the results in prior work, we clarify the distinctions between our setup and previous ones. The first difference is in the order of magnitude of the scores reported in our work. Prior work like Fontaine 2021 (Fontaine & Nikolaidis, 2021) and Fontaine 2023 (Fontaine & Nikolaidis, 2023) reported normalized QD scores, which are QD scores divided by the number of cells in the archive (10,000). Those works also normalized objective values to be 0-100, whereas we normalize to 0-1 to make the objective values more amenable to the discount model. These two changes explain the difference in orders of magnitude between our results and those works: when divided by 10,000 and multiplied by 100, our results for CMA-MAE, MAP-Elites, and MAP-Elites (line) on 2D LP (Sphere), 2D LP (Rastrigin), and Arm Repertoire are nearly identical to those for LP (sphere), LP (Rastrigin), and Arm Repertoire in Fontaine 2023 (our remaining benchmark domains were not included in Fontaine 2023).
@@ B.1 Comparability of Results to Prior Work
Figure 3: Mean and standard error of the mean for QD Score and Coverage of each algorithm in each domain. Standard error may not be visible in some plots.
@@
Figure 4: Mean and standard error of the mean for QD Score and Coverage of each algorithm in each domain. Standard error may not be visible in some plots.
@@
Figure 5: A random subset of images generated by DMS in the TA (MNIST) domain, where desired measures are sampled from the MNIST dataset. The goal in this domain is to arrange triangles to look like the given MNIST digits. Each rendered triangle image is shown to the left of its corresponding MNIST digit.
Figure 6: A random subset of images generated by DMS in the TA (F-MNIST) domain, where desired measures are sampled from the Fashion MNIST dataset. The goal in this domain is to arrange triangles to look like the images of fashion items. Each rendered triangle image is shown to the left of its corresponding fashion item.
Figure 7: Heatmaps of a randomly selected archive produced by each algorithm in domains with 2D measure spaces. Each row contains heatmaps for a single domain. The axes of the heatmaps are the measures, while the color of each cell indicates the objective value. Notably, the heatmaps show how DMS achieves high coverage of the measure space. They also show how DDS achieves good coverage but cannot achieve high objective values since it is a diversity optimization algorithm.
@@
Figure 8: Progression of the archive and discount model in DMS in the 2D LP (Sphere) benchmark. The left heatmap shows the archive, while the right heatmap shows the discount model. To plot the discount model, we computed its output at points in a $200\times 200$ grid in measure space. The discount model heatmap also shows the dataset $\mathcal{D}_{A}$ of points on a given iteration — blue circles indicates points created with solutions from the emitters, and yellow triangles indicate empty points. On Iteration 0, the discount model initializes to output $f_{min}$ everywhere. On Iteration 250, as the emitters begin to populate the archive, the discount model begins to output higher values in areas that have been explored. However, unexplored areas still maintain low values (shown as dark colors) due to the empty points in the dataset. On further iterations, the discount model outputs higher and higher values as the emitters populate the archive further, until it outputs high values nearly everywhere on Iteration 10000.
Table 3: Welch’s one-way ANOVA results in each domain. All $p$-values are less than 0.001.
@@
Table 4: Pairwise comparisons (Games-Howell test) of the QD Score of each algorithm.
@@
Table 5: Pairwise comparisons (Games-Howell test) of the Coverage of each algorithm.
@@
Appendix C Hyperparameters
Table 6: Hyperparameters.
@@ Appendix C Hyperparameters
Table 6 lists the hyperparameters of DMS and all baseline algorithms. All algorithms run for 10,000 iterations. The number of solutions generated and evaluated on each iteration is equal across all algorithms. In DMS, CMA-MAE, and DDS, this number is equivalent to the number of emitters $W$ times the emitter batch size $\lambda$. In the two MAP-Elites algorithms, this number is equivalent to the batch size $\lambda$. For DDS, we use the KDE version (“DDS-KDE”) (Lee et al., 2024). In all domains, the objective is normalized to be between 0 and 1, so the minimum objective $f_{min}$ is set to 0. For the benchmark domains, parameters for the baselines are adapted from prior work (Fontaine & Nikolaidis, 2023; Lee et al., 2024).
@@
Restart Rule. The restart rule refers to the conditions upon which emitters are restarted from solutions in the archive. “No Imp.” refers to restarting when the emitter no longer discovers solutions that are added to the archive (i.e., solutions that improve the archive) (Fontaine et al., 2020). “Basic” refers to restarting only when default CMA-ES (Hansen, 2016) termination rules are met. An integer value $R$ refers to restarting every $R$ iterations (Fontaine & Nikolaidis, 2023). We study the role of the restart rule in DMS in further detail in Appendix D.
@@
C.1 On the Choice of MLP Discount Models
@@
C.2 On Nearest Neighbor Search in Archives
@@
Appendix D Ablation Study
D.1 Archive Learning Rate
@@
Figure 9: Mean and standard error of the mean of QD Score and Coverage when varying the archive learning rate $\alpha$ in DMS in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@
Figure 10: Mean and standard error of the mean of QD Score and Coverage when varying the number of empty points $n_{empty}$ in DMS in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@
Table 7: Mean and standard error of the mean of QD Score and Coverage with a basic restart rule and a restart rule of 100 in the benchmark domains. Highlighted lines indicate results from the main paper in Table 1. Mean over 20 trials.
@@
Iteration 0
Iteration 250
Iteration 10000
Figure 11: Similar to Fig. 8, this figure shows how the archive and discount model in DMS progress across iterations. However, this time, DMS does not train the discount model with any empty points, i.e., $n_{empty}=0$. As a result, the discount model takes on arbitrary values in areas of the measure space that have not been explored yet, as evinced by the high values across the discount model heatmap on Iteration 250 and 10000. Because the discount values are high everywhere, the emitters in DMS mistakenly believe they have explored all areas of the measure space, even though the archive is essentially empty.
D.2 Empty Points
D.3 Restart Rule
The “restart rule” in DMS and CMA-MAE refers to the conditions upon which the emitters restart search from a new emitter in the archive (Algorithm 1, line 1-1). By default, emitters in DMS and CMA-MAE use a “basic” restart rule, which makes the emitter restart when the default termination conditions for CMA-ES (Hansen, 2016) are met, e.g., the area of the search distribution becomes too small. While tuning DMS, we found it helpful in some domains to make the emitters restart on a fixed schedule as introduced in prior work (Fontaine & Nikolaidis, 2023), e.g., restarting every 100 iterations. To better understand the effect of the restart rule in DMS, we thus present an ablation where we run DMS in the benchmark domains with both a “basic” restart rule and a restart rule of 100.
Appendix E Domain Details
E.1 Linear Projection
@@
Figure 12: Irwin-Hall distribution when summing $r$ uniformly random variables with range [5.12,5.12][-5.12,5.12].
@@
E.2 Arm Repertoire
Arm Repertoire (Cully & Demiris, 2017; Vassiliades & Mouret, 2018) considers solutions ${\bm{\theta}}\in\mathbb{R}^{n}$ that represent the $n$ joint angles of a planar robotic arm with $n$ joints and $n$ links of length 1. The measure function, ${\bm{m}}({\bm{\theta}})$, uses forward kinematics to compute the $(x,y)$ position of the end effector — thus, the measure space is two-dimensional. The objective function seeks to minimize the variance between each joint angle:
@@ E.2 Arm Repertoire
E.3 Triangle Arrangement (TA)
This QDDM domain considers arranging triangles to create images, as inspired by prior work (Tian & Ha, 2022). We consider grayscale images in our work, with the triangles all drawn on a black background. Each triangle is parameterized by the following eight values:
@@ E.3 Triangle Arrangement (TA)
E.4 Latent Space Illumination (LSI)
We construct this QDDM domain by adapting prior work that introduced LSI domains with 2D measures (Fontaine & Nikolaidis, 2021; 2023), and prior work that shows how to guide StyleGAN3 to generate images that match text prompts (Herrera-Berg, 2021). In our domain, which we refer to as LSI (Hiker), the goal is to generate face images of hikers in different landscapes, e.g., a hiker who is ready for the mountains.
@@ Appendix F Implementation
Compute Usage. In benchmark domains, we run all 20 trials of each algorithm in parallel. In QDDM domains, we run algorithms serially. For wallclock times, refer to Table 2. In all domains, we train the discount model of DMS on the GPU. The discount model typically requires 300-400MB of GPU memory. In TA (MNIST) and TA (F-MNIST), we accelerate rendering of the triangles on the GPU. In LSI (Hiker), we accelerate StyleGAN3 and CLIP with the GPU. The final experimental results required 100GB of storage.
@@
Appendix G Differentiable Quality Diversity in QDDM Domains
While our work considers a black-box QD setting as defined in Sec. 2, prior work (Fontaine & Nikolaidis, 2021) introduced differentiable quality diversity (DQD), a QD setting where the objective and measure functions are first-order differentiable. Since DQD methods make assumptions that DMS and our other baselines do not, viz., that the objective and measures are differentiable, we have not included DQD methods in our experiments. Nevertheless, here we discuss considerations in applying DQD methods to QDDM domains.
The primary consideration is that DQD methods, such as CMA-MEGA (Fontaine & Nikolaidis, 2021), were designed with the assumption of a high-dimensional solution space that could be reduced to a low-dimensional objective-measure space, under the assumption that the measure space is low-dimensional. To elaborate, during operation, CMA-MEGA and other DQD methods compute a gradient for the objective and for each measure. In low-dimensional measure spaces, e.g., a 2D measure space, this works well because only a few gradients, 1 for the objective and 2 for the measures in this case, need to be computed. In contrast, in QDDM domains, the measure spaces are much higher-dimensional than even the solution space. In LSI (Hiker), the solution space is 512D while the measure space is 196,608D ($256\times 256\times 3$ RGB image). In the TA domains, the solution space is 240D while the measure space is 784D (it is also worth noting that the TA domains are non-differentiable due to the rendering process). As such, running DQD in these domains would mean computing up to 196,608 measure gradients on every iteration, which is computationally intractable. Given this limitation, we believe developing a DQD method that operates in QDDM domains is an exciting avenue for future work.